OpenStudy (anonymous):

Conservation of Energy 1. A large chunk of ice with mass 15.0 kg falls from a roof 8.00 m above the ground. a. Ignoring air resistance, find the kinetic energy of the ice when it is about to reach the ground? b. What is the speed of the ice immediately before it reaches the ground?

OpenStudy (kmullis6):

idk if this means anything (i haven't started this stuff yet) , but i looked it up online and this came up... Only force acting on the ice is gravity which is conservative therefore E_i = E_f P_i + K_i = P_f + K_f Initially, the chunk is stationary so K_i = 0 Finally the height will be approximately at the ground, thus P_f = 0 P_i = K_f K_f = m*g*h_i = 15 [kg] * 9.81 [m/s^2] * 8 [m] = 1177 [J] <------- ANS K_f = 0.5*m*v^2 v = sqrt(2*K_f / m) = 12.53 [m/s] <------- ANS really hope this helps (:

OpenStudy (anonymous):

loll i posted that question on yahoo answers and got that answer haha. didnt really understand it. really appreciate your time though. Thank you so much

OpenStudy (kmullis6):

haha im sorry my bad /: i'm sorry i couldn't help

OpenStudy (anonymous):

nahh i appreciate it